b) Let 







σ =
3 4 5 2 1
1 2 3 4 5
and 







τ =
5 3 2 4 1
1 2 3 4 5
in . S7
Write στ as a product
of disjoint permutations.
Further, is στ even? Why, or why not?
1
Expert's answer
2020-04-13T16:57:39-0400

Given that ,

"\sigma=\begin{bmatrix}
 3&4&5&2&1 \\
 1&2&3&4&5
\end{bmatrix},\gamma =\begin{bmatrix}
 5&3&2&4&1 \\
 1&2&3&4&5
\end{bmatrix} \in S_7"


"\sigma\gamma=\begin{bmatrix}
 3&4&5&2&1 \\
 1&2&3&4&5
\end{bmatrix}\begin{bmatrix}
 5&3&2&4&1\\
 1&2&3&4&5
\end{bmatrix}"


"=\begin{bmatrix}
 1&2&3&4&5 \\
 3&1&4&2&5
\end{bmatrix}" "=(1 \ 3 \ 4 \ 2)" .

"=(1 \ 2)(1 \ 4)(1 \ 3)"

Since we know that every permutation in "S_n,n>1," is a product of 2-cycles. However, the decomposition of a permutation into a product of two cycles is not unique, but the number of 2-cycles is equal in every decomposition,

i.e., if"\, \alpha=\beta_1.\beta_2.......\beta_r \ \ and \ \alpha=\gamma_2.\gamma_2............\gamma_s"

where "\beta's \ and \ the \ \gamma's" are 2-cycles.

Then "r=s" .

Also we know that a permutation can be expressed as a product of an odd number of 2-cycles is called an odd permutation.

Hence by definition, "\sigma\gamma" is an odd permutation.